BIBLIOGRAFIJA PROF. MIRJANE VUKIĆEVIĆ - KARABIN

SPISAK RADOVA

1. RESULTS OF IONOSPHERIC SOUNDING ABOVE BEOGRAD DURING THE SOLAR ECLIPSE OF 15-TH FEBRUARY 1961, 1962, Bulletin scien. Conseil Acad. R.P.F. Yougoslavie, \underline{7}, No.6 (koautor D.Bajić)

2. IONOSPHERIC EFFECTS FOLLOWING DISTANT NUCLEAR DETONATIONS, 1963, Journ. Atmospheric and Terrest. Physics, London, \underline{25}, 212–218.

3. UTICAJ NUKLEARNIH EKSPLOZIJA U ATMOSFERI NA PROSTIRANJE RA-DIO TALASA, 1963, Zbornik radova VIII jugoslovenske konferencije ETAN-a (koautor D.Bajić)

4. HROMOSFERSKE ERUPCIJE PRAĆENE RADIO ZRAČENJEM I NJIHOV UTI-CAJ NA JONOSFERU I GEOMAGNETIZAM, 1963, Bilten društva matematičara SR Srbije, \underline{15}, No. 1–4 (koautor R. Vojvodić)

5. LEVEL OF SOLAR FLUX AT 3000 Mc/s AND ITS RELATION TO OCCURENCE OF SID’s, 1964, Journ. Atm. and Terr. Phys., London, \underline{26}, 1138–1142. (koautor A.P.Mitra)

6. STATISTICAL CHARACTERISTICS OF SUDDEN IONOSPHERIC DISTUR-BANCES, 1965, RPU Sci. Rep. No. 19, New Delhi

7. IZNENADNI JONOSFERSKI POREMEĆAJI – SID I METODE NJIHOVE DE-TEKCIJE I PROUČAVANJA, 1965, PMF – Beograd (doktorska disertacija)

8. POUZDANOST PROGNOZIRANJA MU FREKVENCIJE ZA JONOSFERU IZ-NAD NAŠE ZEMLJE, 1968, Zbornik radova XII jugoslovenske konferencije ETAN–a, 224–232.

9. FLUX OF SOLAR XUV RADIATION INFERRED FROM IONOSPHERIC VARI-ATIONS DURING THE ECLIPSE OF 20-TH MAY 1966, 1969, Publ. of Dept. of Astronomy, Faculty of Sciences, Beograd, No. 1, 3–7.

10. LOW ENERGY PARTICLE EFFECTS OVER BELGRADE AFTER PROTON FLARE OF THE 7–TH JULY 1966, 1970, Publ. of Dept. of Astronomy, Faculty of Sciences, Beograd, No.2, 1–4.

11. ANOMALNA APSORPCIJA RADIO–TALASA U JONOSFERI, 1970, Zbornik radova XIV konferencije ETAN–a, 117–124.

12. THE CRITERIA FOR HOMOLOGOUS RADIO EVENTS ASSOCIATED WITH FLARES, 1971, Publ. of Dept. of Astronomy, Faculty of Sciences, Beograd, No.3 (koautor M. Dimitrijević)

13. RADIO–KOMUNIKACIJE POMOĆU E_{S} SLOJA, 1972, Zbornik radova XVI konferencije ETAN–a

14. ON THE STRUCTURE AND DEVELOPMENT OF THE SUNSPOT GROUPS, 1974, Publ. of Dept. of Astronomy, Faculty of Sciences, Beograd, No.5, 9–17 (koautor P. Razinger)

15. DOGADJAJI NA SUNCU AVGUSTA 1972. g. I NJIHOVE POSLEDICE U ZEMLJINOJ ATMOSFERI, 1975, Zbornik I nacionalne konferencije jugoslovenskih astronoma, Publ. Astr. Obs. Beograd, No.20.

16. O NEKIM NEREŠENIM PROBLEMIMA FIZIKE SUNCA, 1975, Zbornik VI kongresa MFAJ, Novi Sad, 247.

17. SOME RESULTS OF THE PHOTOSPHERIC LARGE SCALE VELOCITY RE-SEARCH FROM BELGRADE OBSERVATIONS, 1977, Solar Physics \underline{52}, 199–200. (koautor A. Kubičela)

18. LIMB EFFECT ALONG THE CENTRAL MERIDIAN OF THE SOLAR DISC, 1977, Publ. of Dept. of Astronomy, Beograd, No.6, 9–15. (koautor A. Kubičela)

19. APPARENT YEARLY PRECESSION OF THE SUN, 1977, Solar Physics \underline{54}, 505– 509. (koautor A. Kubičela)

20. AN INTEGRAL SUPERGRANULAR CONTRIBUTION TO THE GLOBAL SCALE LINE–OF–SIGHT VELOCITY FIELD, 1979, Zbornik radova IV konferencije jugoslovenskih astronoma, Sarajevo. (koautor A. Kubičela)

21. DETAILED TREATMENT OF SYNODIC SOLAR ROTATION, 1981, Publ. of Dept. of Astronomy, Beograd, No.11, 35–42. (koautor A. Kubičela)

22. LINE–OF–SIGHT VELOCITY FIELD OF SYNODIC SOLAR ROTATION, 1982, in \underline{Sun and Planetary System} (eds. W. Fricke and G. Teleki) D. Reidel, Dordrecht – Holland, 73–76 (koautor A. Kubičela)

23. ANOTHER VIEW OF SYNODIC SOLAR ROTATION, 1983, Solar Physics \underline{84}, 389–393 (koautor A. Kubičela)

24. FURTHER BELGRADE RESULTS OF THE LARGE–SCALE PHOTOSPHERIC VELOCITY RESEARCH, 1986, Bull. Astr. Obs. Beograd \underline{136}, 1. (koautori A. Kubičela, I. Vince, Z. Ivanović)

25. SOLAR AND STELLAR ACTIVITY PHENOMENA, 1986, Bull. Astr. Obs. Beograd \underline{136}, 53. (koautor J. Arsenijević)

26. A BELGRADE PROGRAM FOR MONITORING ACTIVITY SENSITIVE SPEC-TRAL LINES OF SUN AS A STAR, 1987, Proc. of the 5th European Meeting on Solar Physics, Titisee, Freiburg (koautori J. Arsenijević, A. Kubičela)

27. ACTIVITY IN HOMOLOGOUS FLARES, 1987, in Transport of Energy and Masses in Solar Atmosphere (eds. Ph. Lemaire and Z. Mouradian), College de France, Paris.

28. BEGINNING OF A STUDY OF LONG–TERM CHANGES OF SELECTED FRAUNHOFER SPECTRAL LINES, 1987, II Workshop Astrophysics in Yugoslavia (ed. M. Dimitrijević) (koautori A. Kubičela, I. Vince)

29. VARIATION IN THE SOLAR CONSTANT, 1988, Publ. of Dept. of Astronomy, Beograd, \underline{16}, 5.

30. ACTIVITY CYCLE VARIATIONS OF SOME PHOTOSPHERIC LINES, 1990, Proc. of the 6th European Meeting on Solar Physics, Debrecen, Hungary, p. 104 (koautori A. Kubičela, J. Skuljan, I. Vince)

31. MAGNETIC FIELDS IN COOL STARS, 1992, Bull. Astron. Belgrade \underline{145}, 59.

32. SOLAR ACTIVITY INFLUENCE ON EQUIVALENT WIDTHS OF SOME PHO-TOSPHERIC LINES, 1992, Bull. Astron. Belgrade \underline{145}, 1. (koautori J. Skuljan, I. Vince, A. Kubičela)

33. DESIGN RATIONAL OF THE SOLAR ULTRAVIOLET NETWORK (SUN), 1992, Proc. of the ESO Conference, part II, p. 995–1004 (eds. J.M.Beckers and F. Merkle)

34. LONG–TERM CHANGES OF EQUIVALENT WIDTHS ON TEN SELECTED FRAUNHOFER LINES, 1992, JOSO Annual Report, p. 74–77 (ed. A.V. Alvensleben)

35. OSCILLATIONS AND WAVES IN THE SUN AND STARS (Introductory lecture), 1993, Proc. of the X Nat. conf. of Yug. astron. Belgrade (eds. M. Dimitrijević and D. Djurović)

36. LONG–TERM CHANGES OF SOME FRAUNHOFER LINES IN THE SUN AS A VARIABLE STAR, 1993, IAU Colloquium 143, D. Reidel, Dordrecht-Holland (koautori I. Vince, A. Kubičela, J. Skuljan, L. Popović)

UDŽBENICI

1. B.M. Ševarlić, M. Vukićević –Karabin i S. Sadžakov ASTRONOMIJA, za IV razred gimnazije prirodno-matematičkog smera 1975, Zavod za udžbenike i nastavna sredstva, Beograd

2. M. Vukićević – Karabin ASTROFIZIKA, za IV razred usmerenog obrazovanja, 1988, Naučna knjiga, Beograd

3. M. Vukićević – Karabin TEORIJSKA ASTROFIZIKA (univerzitetski udžbenik) 1994, Naučna knjiga, Beograd